Presented below are selected transactions on the books of Culver Corporation. May 1, 2020 Bonds payable with a par value of $861,600, which are dated January 1, 2020, are sold at 106 plus accrued interest. They are coupon bonds, bear interest at 11% (payable annually at January 1), and mature January 1, 2030. (Use interest expense account for accrued interest.) Dec. 31 Adjusting entries are made to record the accrued interest on the bonds, and the amortization of the proper amount of premium. (Use straight-line amortization.) Jan. 1, 2021 Interest on the bonds is paid. April 1 Bonds with par value of $344,640 are called at 102 plus accrued interest, and redeemed. (Bond premium is to be amortized only at the end of each year.) Dec. 31 Adjusting entries are made to record the accrued interest on the bonds, and the proper amount of premium amortized. Prepare journal entries for the transactions above.
Answer #1

B C D 1 Journal Entry 2 Date Particulars debit credit 944888 4 01-May-17 Cash Bonds Payable Premium on Bonds Payable Interest21 22 23 01-Apr-18 Bonds Payable Premium on Bonds Payable Interest Expense Cash Gain on Redemption of Bonds 344640 18718 94781 Journal Entry 2 Date Particulars debit credit 4. 42856 =861600*106%+(861600*11%*4/12) =C4 covou Cash Bonds Payable Premium22 43191 Bonds Payable Premium on Bonds Payable Interest Expense Cash Gain on Redemption of Bonds 344640 |=51696 *(344640 / 8